OpenStudy (anonymous):

Please Help! 3(5)^(2x-3) = 6??

OpenStudy (mertsj):

1. Divide both sides by 3 2. Take the log of both sides 3. Divide both sides by log 5 4. Add 3 to both sides 5. Divide both sides by 2

OpenStudy (anonymous):

I don't have a calculator with me @Mertsj

OpenStudy (mertsj):

use an online calculator.

OpenStudy (anonymous):

@Mertsj Ok, so when I plug it into the calculator it will look something like this then right? log5^(2x03)/log2

OpenStudy (mertsj):

(2x-3)log 5 = log 2

OpenStudy (mertsj):

You need to learn this:\[\log 5^{2x-3}=\log 2\] is the same as: \[(2x-3)\log 5 = \log 2\]

OpenStudy (anonymous):

@Mertsj I got 1.04 rounded up

OpenStudy (mertsj):

log2 divided by log 5 = .4307

OpenStudy (mertsj):

So \[2x-3=.4307\]

OpenStudy (mertsj):

Now add 3 to both sides and then divide both sides by 2